A gold coin is 84 years older than a silver coin. Twenty years ago, the gold coin was three times as old as the silver coin was then. Find the present ages of the two coins.
Answer by stanbon(75887)

A gold coin is 84 years older than a silver coin. Twenty years ago, the gold coin was three times as old as the silver coin was then. Find the present ages of the two coins.
------
Equations:
g = s + 84
----------------
g-20 = 3(s-12)
-------------------
Substitute and solve for "s":
s+84-20 = 3s-36
2s = 100
s = 50 yrs. (age of the silver coin)
------
Solve for "g":
g = s + 84
g = 50+84 = 134 yrs. (age of the gold coin)
=====================
Cheers,
Stan H.
=====================